# Exploratory Randomization for Discrete-Time Linear Exponential Quadratic Gaussian (LEQG) Problem

### v1

@@ 3.3 Solving the Risk-Neutral Penalized Game Problem
$-\left(\mathfrak{A}^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{C}_{t+1}^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}+\left(\mathfrak{A}^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$ (3.38)
$\mathfrak{k}_{t+1}=$ $-\frac{1}{2}\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$
$+\frac{1}{2}\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{C}_{t+1}^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{a}^{(1)}_{t+1}+\frac{1}{2}\left(\mathfrak{a}^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$ (3.39)
@@
$F(u^{*},\gamma^{*},\eta^{*})=$ $\frac{1}{2}x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}K^{(1)}_{t+1}x_{t}+x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}k^{(1)}_{t+1}+\frac{1}{2}\left(k^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}k^{(1)}_{t+1}$
$-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}x_{t}-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(k^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}x_{t}$
@@
$+\frac{1}{2}x_{t}^{\prime}\left(K^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{B}^{(2)}_{t+1}(\mathfrak{G}_{t+1})^{-1}K^{(2)}_{t+1}x_{t}+x_{t}^{\prime}\left(K^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{B}^{(2)}_{t+1}(\mathfrak{G}_{t+1})^{-1}k^{(2)}_{t+1}$
@@
$-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}K^{(2)}_{t+1}x_{t}-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}k^{(2)}_{t+1}$
$-\left(k^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}K^{(2)}_{t+1}x_{t}-\left(k^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}k^{(2)}_{t+1}$ (3.41)
@@
    $P_{t}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}-2\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{A}^{(2)}_{t+1}$
    $+2\left(\theta M+\frac{1}{2}A^{\prime}P_{t+1}A\right),\quad P_{T}=M_{T},$ (3.43)
@@
    $p_{t}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$
@@
    $+A^{\prime}P_{t+1}a+\theta m+A^{\prime}p_{t+1},\quad p_{T}=m_{T},$ (3.44)
@@
    $r_{t}=$ $-\frac{1}{2}\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$
    $+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi N_{t}\right),$
    $r_{T}=0.$ (3.45)
Furthermore, the candidate optimal controls are given by the stationary point $(u^{*},\gamma^{*},\eta^{*})$ of $F(\bar{u},\bar{\gamma},\bar{\eta})$ according to Proposition 3.3 (formulas (3.3),(3.3), (3.3))
@@
$V_{t}(x_{t})=$ $F(u_{t}^{*},\gamma_{t}^{*},\eta_{t}^{*})+x_{t}^{\prime}\left(\theta M+\frac{1}{2}A^{\prime}P_{t+1}A\right)x_{t}+x_{t}^{\prime}\left(A^{\prime}P_{t+1}a+\theta m+A^{\prime}p_{t+1}\right)$
$+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi N_{t}\right)$
@@
$+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi N_{t}\right).$ (3.46)
@@
$\inf_{\bar{u}}I(u;T,\theta)=\exp\left\{\frac{1}{2}x_{t}^{\prime}P_{t}x_{t}+x_{t}^{\prime}p_{t}+r_{t}\right\}$ (3.47)
@@
$\sup_{\bar{u}}J(u;T,\theta)=-\frac{1}{\theta}\left(\frac{1}{2}x_{t}^{\prime}P_{t}x_{t}+x_{t}^{\prime}p_{t}+r_{t}\right)$ (3.48)
with $P_{t}$, $p_{t}$, and $r_{t}$ given by (i),(ii), and (iii), and optimal controls given by the stationary point $(u^{*},\gamma^{*},\eta^{*})$ according to Proposition 3.3.
@@
and (3.48) follows from the definition of $I$ at (2.4). ∎
@@ 3.4 Sufficient conditions for the existence of a saddle point
According to the proof of Proposition 3.3, the search for a saddle point for $V_{t}(x_{t})$ reduces to the search of a saddle point of $F(\bar{u},\bar{\gamma},\bar{\eta})$ in the sense of (3.29). It means that, at the stationary point $(u^{*},\eta^{*},\gamma*)$, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ must have a minimum in $\bar{u}$ for all values of $\bar{\nu}=(\bar{\gamma}^{\prime},\bar{\eta}^{\prime})^{\prime}$. A sufficient condition for this to happen is that, at $(u^{*},\eta^{*},\gamma*)$, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ is convex in $\bar{u}$. Analogously, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ must have a maximum in $\nu^{*}=({\gamma^{*}}^{\prime},{\eta^{*}}^{\prime})^{\prime}$ for all values of $\bar{u}$. A sufficient condition is that, at $(u^{*},\eta^{*},\gamma^{*})$, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ is concave in $\bar{\nu}$. We do not need further global convexity/concavity properties of $F(\bar{u},\bar{\gamma},\bar{\eta})$. Therefore, we can apply the usual second-order sufficiency condition separately to $\bar{u}$ and $\bar{\nu}$, namely
@@
$-H^{\bar{\gamma},\bar{\eta}}_{t}=-\begin{pmatrix}\mathfrak{B}^{(2)}_{t+1}&\mathfrak{C}_{t+1}^{\prime}\\ F\mathfrak{C}_{t+1}&\mathfrak{B}^{(3)}_{t+1}\end{pmatrix}>0,$ (3.54)
@@
Notice that conditions such as (3.53) are standard for both LQG and LEQG problems and that condition (3.54) results from the energy-entropy duality. According to the Sylvester conditions, we can express the negative definiteness condition (3.54) as the following three simultaneous conditions
@@
The second condition in (3.55) is the second-order condition we would have if we maximized for $\bar{\gamma}$ only, ignoring $\bar{\eta}$. As $\Lambda_{t}$ and $P_{t+1}$ are both deterministic, this condition can be checked in advance, even if the recursion for $P_{t}$ depends on $\Lambda_{t+1}$ via the terms $\mathfrak{B}^{(2)}_{t+1}$ and $\mathfrak{G}_{t+1}$, defined respectively at (3.3) and (3.3). This second condition in (3.55) is also reminiscent of the risk-resistance condition found in the LEQG literature (Shaiju & Petersen,, 2008). This was to be expected. The energy-entropy duality expresses the original LEQG problem as an LQG problem. Hence, any condition necessary for the resolution of the LEQG problem should also apply in some way to the equivalent LQG game. Such a condition will manifest through the penalizing control $\gamma$, which is the driving term of the change of measure in the energy-entropy duality.
Analogously, the last condition in (3.55) is the second-order condition if we maximized for $\bar{\eta}$ and ignored $\bar{\gamma}$. This condition sets a bound on the covariance of exploration $\Xi_{t},$ based on the parameters of the problem, $\theta,N_{t},B$, and the quadratic coefficient $P_{t+1}$. As $\Xi_{t}$ and $P_{t+1}$ are deterministic, an appropriate value can be selected in advance for $\Xi_{t}$.

### v2

@@ 3.3 Solving the Risk-Neutral Penalized Game Problem
$\mathfrak{k}_{t+1}=$ $-\frac{1}{2}\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$ (3.39)
@@
$F(u^{*},\gamma^{*},\eta^{*})$ (3.41)
$=$ $\frac{1}{2}x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}K^{(1)}_{t+1}x_{t}+x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}k^{(1)}_{t+1}$
$+\frac{1}{2}\left(k^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}k^{(1)}_{t+1}-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}x_{t}$
$-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(k^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}x_{t}$
@@
$+\frac{1}{2}x_{t}^{\prime}\left(K^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{B}^{(2)}_{t+1}(\mathfrak{G}_{t+1})^{-1}K^{(2)}_{t+1}x_{t}$
$+x_{t}^{\prime}\left(K^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{B}^{(2)}_{t+1}(\mathfrak{G}_{t+1})^{-1}k^{(2)}_{t+1}$
@@
$-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}K^{(2)}_{t+1}x_{t}$
$-x_{t}^{\prime}\left(K^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}k^{(2)}_{t+1}$
$-\left(k^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}K^{(2)}_{t+1}x_{t}$
$-\left(k^{(1)}_{t+1}\right)^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{C}_{t+1}\left(\mathfrak{G}_{t+1}\right)^{-1}k^{(2)}_{t+1}$
@@
    $P_{t}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{A}^{(1)}_{t+1}$ (3.43)
    $-2\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{A}^{(2)}_{t+1}$
    $+\left(\mathfrak{A}^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{A}^{(2)}_{t+1}+2\left(\theta M+\frac{1}{2}A^{\prime}P_{t+1}A\right),$
    $P_{T}$ $=M_{T},$
@@
    $p_{t}=$ $-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}-\left(\mathfrak{A}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$ (3.44)
@@
    $+A^{\prime}P_{t+1}a+\theta m+A^{\prime}p_{t+1},$
    $p_{T}=$ $m_{T},$
@@
    $r_{t}=$ $-\frac{1}{2}\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{a}^{(1)}_{t+1}$ (3.45)
    $-\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$
    $+\frac{1}{2}\left(\mathfrak{a}^{(1)}_{t+1}\right)^{\prime}(\mathfrak{B}^{(1)}_{t+1})^{-1}\mathfrak{C}_{t+1}(\mathfrak{G}_{t+1})^{-1}\mathfrak{C}_{t+1}^{\prime}\left(\mathfrak{B}^{(1)}_{t+1}\right)^{-1}\mathfrak{a}^{(1)}_{t+1}$
    $+\frac{1}{2}\left(\mathfrak{a}^{(2)}_{t+1}\right)^{\prime}(\mathfrak{G}_{t+1})^{-1}\mathfrak{a}^{(2)}_{t+1}$
    $+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}$
    $+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi_{t}N_{t}\right),$
    $r_{T}=0.$
Furthermore, the candidate optimal controls are given by the stationary point $(u^{*},\gamma^{*},\eta^{*})$ of $F(\bar{u},\bar{\gamma},\bar{\eta})$ according to Proposition 3.3 (formulas $(3.3),$ (3.3))
@@
$V_{t}(x_{t})$ (3.46)
$=$ $F(u_{t}^{*},\gamma_{t}^{*},\eta_{t}^{*})+x_{t}^{\prime}\left(\theta M+\frac{1}{2}A^{\prime}P_{t+1}A\right)x_{t}+x_{t}^{\prime}\left(A^{\prime}P_{t+1}a+\theta m+A^{\prime}p_{t+1}\right)$
$+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi_{t}N_{t}\right)$
@@
$+\frac{1}{2}\text{tr}\left(B^{\prime}\Xi_{t}BP_{t+1}\right)+\frac{1}{2}\text{tr}\left(\Lambda_{t}P_{t+1}\right)+r_{t+1}+\frac{1}{2}a^{\prime}P_{t+1}a+a^{\prime}p_{t+1}+\theta\text{tr}\left(\Xi_{t}N_{t}\right).$
@@
$\inf_{\bar{u}}I(\bar{u};T,\theta)=\exp\left\{\frac{1}{2}x_{t}^{\prime}P_{t}x_{t}+x_{t}^{\prime}p_{t}+r_{t}\right\}$ (3.47)
@@
$\sup_{\bar{u}}J(\bar{u};T,\theta)=-\frac{1}{\theta}\left(\frac{1}{2}x_{t}^{\prime}P_{t}x_{t}+x_{t}^{\prime}p_{t}+r_{t}\right)$ (3.48)
with $P_{t}$, $p_{t}$, and $r_{t}$ given by (3.43),(3.44), and (3.45), and optimal controls given by the stationary point $(u^{*},\gamma^{*},\eta^{*})$ according to Proposition 3.3.
@@
and (3.48) follows from the definition of II at (2.4). ∎
@@ 3.4 Sufficient conditions for the existence of a saddle point
According to the proof of Proposition 3.3, the search for a saddle point for $V_{t}(x_{t})$ reduces to the search of a saddle point of $F(\bar{u},\bar{\gamma},\bar{\eta})$ in the sense of (3.29). It means that, at the stationary point $(u^{*},\eta^{*},\gamma*)$, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ must have a minimum in $\bar{u}$ for all values of $\bar{\nu}=(\bar{\gamma}^{\prime},\bar{\eta}_{t}^{\prime})^{\prime}$. A sufficient condition for this to happen is that, at $(u^{*},\eta^{*},\gamma*)$, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ is convex in $\bar{u}$. Analogously, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ must have a maximum in $\nu^{*}=({\gamma^{*}}^{\prime},{\eta^{*}}^{\prime})^{\prime}$ for all values of $\bar{u}$. A sufficient condition is that, at $(u^{*},\eta^{*},\gamma^{*})$, the function $F(\bar{u},\bar{\gamma},\bar{\eta})$ is concave in $\bar{\nu}$. We do not need further global convexity/concavity properties of $F(\bar{u},\bar{\gamma},\bar{\eta})$. Therefore, we can apply the usual second-order sufficiency condition separately to $\bar{u}$ and $\bar{\nu}$, namely
@@
$H^{\bar{\gamma},\bar{\eta}}_{t}=-\begin{pmatrix}\mathfrak{B}^{(2)}_{t+1}&\mathfrak{C}_{t+1}^{\prime}\\ \mathfrak{C}_{t+1}&\mathfrak{B}^{(3)}_{t+1}\end{pmatrix}>0,$ (3.54)
@@
Notice that conditions such as (3.53) are standard for both LQG and LEQG problems, and that condition (3.54) results from the energy-entropy duality. According to Sylvester’s criterion, we can express the negative definiteness condition (3.54) as the following three simultaneous conditions
@@
The second condition in (3.55) is the second-order condition we would have if we maximized for $\bar{\gamma}$ only, ignoring $\bar{\eta}$. As $\Lambda_{t}$ and $P_{t+1}$ are both deterministic, this condition can be checked in advance, even if the recursion for $P_{t}$ depends on $\Lambda_{t+1}$ via the terms $\mathfrak{B}^{(2)}_{t+1}$ and $\mathfrak{G}_{t+1}$, defined respectively at (3.19) and (3.20). This second condition in (3.55) is also reminiscent of the risk-resistance condition found in the LEQG literature (Shaiju and Petersen, 2008). This was to be expected. The energy-entropy duality expresses the original LEQG problem as an LQG problem. Hence, any condition necessary for the resolution of the LEQG problem should also apply in some way to the equivalent LQG game. Such a condition will manifest through the penalizing control $\gamma$, which is the driving term of the change of measure in the energy-entropy duality.
Analogously, the last condition in (3.55) is the second-order condition if we maximize for $\bar{\eta}$ and ignore $\bar{\gamma}$. This condition sets a bound on the covariance of exploration $\Xi_{t},$ based on the parameters of the problem, $\theta,N_{t},B$, and the quadratic coefficient $P_{t+1}$. As $\Xi_{t}$ and $P_{t+1}$ are deterministic, an appropriate value can be selected in advance for $\Xi_{t}$.